OpenStudy (anonymous):

For the function y = f(x), what is the ordered pair for the point on the graph when x = p + 1? (3 points) Question 1 options: 1) (x, f(p)) 2) (x, p + 1) 3) (p + 1, f(p + 1)) 4) (p + 1, f(p))

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

if x=p+1 and y=f(x), then you substitute x into (x,y)... So, the answer is 3.